Fermi National Laboratory


Accelerator Update for September 28, 2005 - September 30, 2005

Wednesday September 28
The day shift began with Operations monitoring a small Recycler stash, with the Tevatron (TeV) recovering from a quench, with the Antiproton Source (Pbar) stacking, with the Collider Detector at Fermilab (CDF) and D-Zero (D0) waiting for the TeV recovery, and with the MiniBooNE, NuMI, and the Meson MCenter experiments taking beam.

At 9:15 AM, a cryo system technician reported a helium leak on a TeV sector A4 transfer line near the A4 wet engine. The TeV had just started a QBS ramp test when the leak was reported. The test was put on hold.

Operators prepared CDF and D0 for access to their collision halls at 9:54 AM.

The Recycler began stashing antiprotons at 10:27 AM. The stash ended and Pbar resumed stacking at 10:53 AM.

Operations reported that the TeV QBS test was completed by 11:24 AM. CDF came out of access.

At 11:13 AM, Operations reported a Pbar vacuum problem that cascaded from AP1 to AP2 and through to AP3.

Operators investigated and discovered a vacuum leak in the AP1 transfer line at 11:16 AM.

Operators prepared the Pbar transfer tunnels for access and experts began their access at 11:49 AM.

D0 came out of access around 11:50 AM.

Operations put the TeV through a dry squeeze at 12:33 PM.

A Switchyard expert began tuning beam to the Meson MTest experiment around 1:10 PM.

A TeV expert began a study at 2:11 PM.

A Pbar expert reported at 2:33 PM that they found a vacuum leak in a small dipole (EB6). They plan to remove the beam pipe, repair the leak, and then replace it.

MTest experimenters reported at 3:56 PM that they were done with beam for the day.

Operations put the TeV into shot setup at 8:47 PM.

Operations established a Recycler-only store 4411 at 10:43 PM with an initial luminosity of 63.63E30. MCenter resumed taking beam.

At 11:33 PM, Operations reported a NuMI Horn trip. It had tripped off a few times throughout the evening shift. Operations contacted NuMI experts.

Thursday September 29
The midnight shift began with Operations monitoring store 4411, with Pbar off due to vacuum repairs, with NuMI off due to Horn problems, and with MiniBooNE and MCenter taking beam.

A NuMI expert got the Horn running and NuMI resumed taking beam at 2:08 AM. Unfortunately, the Horn tripped off only seven minutes later. It reset okay, but NuMI experts had to resume their Horn investigations.

At 4:12 AM, Operation reported that the NuMI beam permit had tripped off three times due to a target helium flow high limit alarm. Operations called a NuMI expert back in.

At 4:32 AM, the NuMI Horn tripped off for the first of two times in eight minutes. It reset each time. The NuMI expert reported that he thought the trips were due to a faulty SCR.

Main Injector (MI) experts began a couple of studies at 9:04 AM, beginning with an RF study and then followed by a slip stacking study.

Operations established beam to the MTest experiment at around 9:16 AM.

At 9:51 AM, experts began setting up the MI RF stations for a study that inhibited all downstream beam for about thirty minutes while the stations were being prepared.

MTest experimenters reported at 9:59 AM that they were done with the beam for the day.

MI resumed accelerating beam at 10:21 AM.

At 10:38 AM, CDF called and said that beam movement had caused them to turn off their detector. A TeV expert tuned the beam and CDF was able to turn back on at 11:25 AM.

The MI RF study ended at 12:19 PM. Operations inhibited beam while the experts put the RF stations back to their normal operating mode. Normal beam resumed at 12:39 PM.

Pbar experts began a two-hour stacktail study at 1:47 PM.

Pbar experts reported at 3:23 PM that all welding was done and operators interlocked the AP1 enclosure. Experts expected it to take about three hours before the vacuum was back to normal.

Operations established beam to the Accumulator at 7:19 PM.

Friday September 30
The midnight shift began with Operations monitoring store 4411, with Pbar stacking, and with MiniBooNE and NuMI taking beam.

Operations began a TeV end-of-store study around 12:06 AM.

The Recycler began preparing to stash at 1:08 AM. The stash ended at 1:56 AM.

Pbar resumed stacking at 2:12 AM.

Operations terminated store 4411 at 5:03 AM. Operators prepared the CDF collision hall for access.

The Plans for Friday and the Weekend
The plans for today are to allow TeV experts to conduct studies. The next shot is scheduled for after the TeV studies, which should occur before noon.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 9/30/2005   email Fermilab